CNNBC recently reported that the mean annual cost of auto insurance is 984 dollars. Assume the standard deviation is 275 dollars. You take a simple random sample of 79 auto insurance policies. (Do not use tables unless directed to do so.)

Find the probability that a single randomly selected value is at least 999 dollars.
P(X > 999) =

Find the probability that a sample of size n=79n=79 is randomly selected with a mean that is at least 999 dollars.
P(¯xx¯ > 999) =

Enter your answers as numbers accurate to 4 decimal places.
Answer #1

Solution :

Given that ,

mean = \mu = 984

standard deviation = \sigma = 275

P(x < 999 ) = P[(x - \mu ) / \sigma < (999 -984) / 275 ]

= P(z < 0.05 )

Using z table,

= 0.5199

   n = 79

\mu\bar x =  \mu = 984

\sigma\bar x = \sigma / \sqrt n = 275 / \sqrt 79= 30.94

P(\bar x < 999 ) = P((\bar x - \mu \bar x ) / \sigma \bar x < (999 - 984) 30.94/ )

= P(z < 0.48 )

Using z table

= 0.6844